# Generative Frame Sampler for Long Video Understanding

### v1

@@ Abstract
[…]
@@
Linli Yao1, Haoning Wu1, Kun Ouyang1, Yuanxing Zhang1, Caiming Xiong2, Bei Chen3, Xu Sun1thanks: Corresponding authors., Junnan Li211footnotemark: 1, 1Peking University, 2Salesforce Research, 3Independent Researcher
@@ 1 Introduction
To summarize, our main contributions are threefold: 1) We propose GenS, a novel generative frame sampler that leverages VideoLLMs to identify question-aware relevant frames. It serves as a plug-and-play sampler that enhances input frames for VideoQA Assistants. 2) We introduce GenS-Video-150K, a large-scale video instruction dataset that densely annotates relevant frames with fine-grained confidence scores across diverse video questions. 3) Through extensive experiments, we demonstrate that GenS significantly enhances the performance of both open-source (Aria-25B, LLaVA-Video-7B/72B) and proprietary (GPT-4o and Gemini) VideoQA Assistants. Notably, when equipped with GenS, LLaVA-Video-72B achieves state-of-the-art performance with accuracy scores of 77.0 on MLVU and 66.8 on LongVideoBench, while Aria attains 39.2 on HourVideo with averaging 45.7 minutes video duration. We will release all models and datasets to facilitate future research.
@@ 2.1 GenS-Video-150K Dataset Collection
We collect 150K data samples from our four-stage pipeline. Each video has an average duration of 647.5 seconds, with 129.5 captioned frames. Among these frames, 20% on average are annotated as relevant with fine-grained confidence scores, providing dense supervision.
@@ 2.3 Training and Inference Paradigm
Table 1: Performance on LongVideoBench (Wu et al., 2024a) and MLVU (Zhou et al., 2024) benchmarks using multiple-choice accuracy metrics. V-Centric denotes a vision-centric subset containing questions that explicitly require video understanding rather than language-only reasoning, while filtering short videos. Frames N/M indicates input N frames for LongVideoBench and M frames for MLVU separately. Using GenS, we select the K most relevant frames (K <= max frame number of VideoQA models) and report the average number of input frames.
@@ 3.1 Experimental Settings
[…]
@@ 4.1 Long-form Video Understanding
4.2 VideoLLM with Retrieval-Augmented Generation
To enhance video-language interaction, recent works have equipped VideoLLM assistants with Retrieval-Augmented Generation (RAG). Unlike text-based retrieval methods, i.e., Video-RAG Luo et al. (2024), Q-ViD Romero and Solorio (2024), and R2A Pan et al. (2023), we propose a visual-centric approach that directly retrieves relevant frames. Compared to CLIP-based retrieval Wang et al. (2024b); Arefeen et al. (2024); Wang et al. (2024c); Xu et al. (2024), our method built on a VideoLLM excels at capturing long-range temporal perception and complex language understanding. While similar frame samplers Yu et al. (2024); Sun et al. (2025) are limited to sparse sampling (e.g., 8 from 128 frames), our approach can efficiently retrieve thousands of frames with adaptive sampling rates, substantially enhancing long-context video assistants on hours-long video perception.
@@ A.4 Frame Sampling Baseline Implementation
You are an advanced AI visual assistant tasked with assessing frame relevance for question answering. Please retrieve the video frames relevant to the question (maybe with options) to answer it correctly, output the frame timestamp, exactly in format [XX:XX], [XX:XX, XX:XX], or [XX:XX, XX:XX, XX:XX], etc. If no matching frames are found, output [None].

### v2

@@ Abstract
Despite recent advances in Video Large Language Models (VideoLLMs), effectively understanding long-form videos remains a significant challenge. Perceiving lengthy videos containing thousands of frames poses substantial computational burden. To mitigate this issue, this paper introduces Generative Frame Sampler (GenS), a plug-and-play module integrated with VideoLLMs to facilitate efficient lengthy video perception. Built upon a lightweight VideoLLM, GenS leverages its inherent vision-language capabilities to identify question-relevant frames. To facilitate effective retrieval, we construct GenS-Video-150K, a large-scale video instruction dataset with dense frame relevance annotations. Extensive experiments demonstrate that GenS consistently boosts the performance of various VideoLLMs, including open-source models (Qwen2-VL-7B, Aria-25B, VILA-40B, LLaVA-Video-7B/72B) and proprietary assistants (GPT-4o, Gemini). When equipped with GenS, open-source VideoLLMs achieve impressive state-of-the-art results on long-form video benchmarks: LLaVA-Video-72B reaches 66.8 (+4.3) on LongVideoBench and 77.0 (+2.7) on MLVU, while Aria obtains 39.2 on HourVideo surpassing the Gemini-1.5-pro by 1.9 points. We release the code, dataset and models at https://generative-sampler.github.io.
@@
Linli Yao1, Haoning Wu2, Kun Ouyang1, Yuanxing Zhang2, Caiming Xiong3, Bei Chen4, Xu Sun1thanks: Corresponding authors., Junnan Li311footnotemark: 1, 1National Key Laboratory for Multimedia Information Processing, School of Computer Science, Peking University 2Peking University, 3Salesforce Research, 4Independent Researcher {linliyao,kunouyang}@stu.pku.edu.cn{longo,xusun}@pku.edu.cn {realtimothyhwu,beibei1019}@gmail.com{cxiong,junnan.li}@salesforce.com
@@ 1 Introduction
To summarize, our main contributions are threefold: 1) We propose GenS, a novel generative frame sampler that leverages VideoLLMs to identify question-aware relevant frames. It serves as a plug-and-play sampler that enhances input frames for VideoQA Assistants. 2) We introduce GenS-Video-150K, a large-scale video instruction dataset that densely annotates relevant frames with fine-grained confidence scores across diverse video questions. 3) Through extensive experiments, we demonstrate that GenS significantly enhances the performance of both open-source (Qwen2VL, Aria, VILA-v1.5, LLaVA-Video) and proprietary (GPT-4o and Gemini-1.5-pro) VideoQA Assistants. Notably, when equipped with GenS, LLaVA-Video-72B achieves state-of-the-art performance with accuracy scores of 77.0 on MLVU and 66.8 on LongVideoBench, while Gemini1.5-pro attains 40.7 on HourVideo with averaging 45.7 minutes video duration.
@@ 2.1 GenS-Video-150K Dataset Collection
Our four-stage pipeline yields 150K data samples, with representative examples shown in Appendix Figure 4. Table 1 provides a comprehensive overview of our dataset statistics. Each video has an average duration of 647.5 seconds (10.8 minutes) and contains approximately 129.5 captioned frames. Of these frames, an average of 20% are annotated as relevant with fine-grained confidence scores ranging from 0 to 5 (where 0 indicates non-relevant and 5 indicates most relevant), thereby providing dense supervision.
Table 1: Statistics of GenS-Video-150K dataset.
Property Value Format {video, question, answer, scored relevant frames} Total Samples 150K Avg Video Duration 647.5 seconds (10.8 minutes) Avg Task Number 12 Relevant Frame Rate around 20% Relevance Scores 0-5 (0: non-relevant, 5: most relevant)
@@ 2.3 Training and Inference Paradigm
Table 2: Performance on LongVideoBench (Wu et al., 2024a) and MLVU (Zhou et al., 2024) benchmarks using multiple-choice accuracy metrics. V-Centric denotes a vision-centric subset containing questions that explicitly require video understanding rather than language-only reasoning, while filtering short videos. Sampled Frames N/M indicates sampled N frames for LongVideoBench and M frames for MLVU separately. Using GenS, we select the K most relevant frames (K <= max frame number of VideoQA models) and report the average number of input frames.
@@ 3.1 Experimental Settings
Training Dataset. We utilize timestamp-output tasks from E.T.Instruct 164K Liu et al. , extracting 75K base training samples (denoted as E.T. Instruct-75K). To enhance the density of grounded frame annotations, we post-process these samples through timestamp label aggregation and textual query concatenation within each video, yielding 41K samples (denoted as $\text{E.T. Instruct-41K}_{\text{agg.}}$). The final training data for GenS combines this aggregated E.T. Instruct dataset with GenS-Video-150K.
@@ 3.5 Analysis
4 Extension Applications
Coarse-to-Fine Hybrid Sampling. We propose a coarse-to-fine hybrid approach that combines a lightweight CLIP Radford et al. (2021) sampler with our GenS to improve sampling efficiency for extremely long videos. Specifically, we first adopt CLIP to densely sample frames from the 1 fps candidate pool and return top 256 most relevant frames, then apply GenS to re-sample the most informative frames within a single 256-frame temporal window.
Table 8 demonstrates that this hybrid approach consistently outperforms both uniform sampling and standalone CLIP sampling across various frame count constraints. We also provide the GenS sampled at 1fps as an upper bound for the frame sampling efficiency-performance tradeoff. This demonstrates that a simple hybrid approach can effectively improve the efficiency of the GenS sampler while maintaining competitive performance, which provides a more practical solution for real-world applications.
#Sampled Frames $\leq$10 frm $\leq$20 frm $\leq$30 frm $\leq$40 frm $\leq$50 frm Uniform 47.04 50.00 53.75 51.98 53.95 CLIP-L-224px 46.44 50.00 53.16 54.55 54.74 CLIP-L-224px + GenS 53.16(+6.12) 56.13(+6.13) 60.47(+6.72) 61.26(+6.71) 60.86(+6.12) GenS (Upper Bound) 56.32 58.89 65.61 65.61 64.82
[…]
#Sampled Frames $\leq$10 frm $\leq$20 frm $\leq$30 frm $\leq$40 frm $\leq$50 frm Uniform 47.04 50.00 53.75 51.98 53.95 InternVL-14B-224px 47.23 49.01 50.99 54.35 54.55 CLIP-L-224px 46.44 50.00 53.16 54.55 54.74 GenSQwen2.5VL-3B 54.74(+7.51) 55.93(+5.93) 57.11(+3.36) 59.68(+5.13) 58.69(+3.95)
[…]
GenS Implementation on Qwen2.5VL-3B. Our design of generative frame sampling is not limited to a specific VideoLLM (e.g., Aria) as the base model. To verify the generalizability of our approach, we implemented GenS on Qwen2.5VL-3B using low-resolution inputs ($\times112$ pixels) for frame sampling.
Results in Table 9 demonstrate that GenS based on Qwen2.5VL-3B achieves remarkable performance compared to both uniform sampling and CLIP-based samplers. The model shows consistent improvements across all frame count configurations, with gains of up to 7.51 points when using just 10 frames. The successful adaptation of GenS to the distinctly different Qwen2.5VL-3B architecture validates the broad generalizability of our approach. Our method can be integrated with various advanced VideoLLMs without requiring architectural modifications, enabling it to leverage ongoing advancements in the field.
@@ 5.1 Long-form Video Understanding
5.2 VideoLLMs with Retrieval-Augmented Generation
To enhance video-language interaction, recent works have equipped VideoLLM assistants with Retrieval-Augmented Generation (RAG). Unlike text-based retrieval methods, i.e., Video-RAG Luo et al. (2024), Q-ViD Romero and Solorio (2024), and R2A Pan et al. (2023), we propose a visual-centric approach that directly retrieves relevant frames. Compared to CLIP-based retrieval Wang et al. (2024b); Arefeen et al. (2024); Wang et al. (2024c); Xu et al. (2024); Tang et al. (2025), our method built on a VideoLLM excels at capturing long-range temporal perception and complex language understanding. While similar frame samplers Yu et al. (2024); Sun et al. (2025) are limited to sparse sampling (e.g., 8 from 128 frames), our approach can efficiently retrieve thousands of frames with adaptive sampling rates, substantially enhancing long-context video assistants on hours-long video perception.
@@ 6 Conclusion
Acknowledgements
This research was partially supported by the National Natural Science Foundation of China under Grant No. 92470205 and No. 62176002. Xu Sun and Junnan Li are the corresponding authors.
@@ A.4 Frame Sampling Baseline Implementation
You are an advanced AI visual assistant tasked with assessing frame relevance for question answering. Please retrieve the video frames relevant to the question (maybe with options) to answer it correctly, output the frame timestamp, exactly in format [$X$:$X$], [$X$:$X$, $X$:$X$], or [$X$:$X$, $X$:$X$, $X$:$X$], etc. If no matching frames are found, output [None].
@@ A.5 Visualization Cases
Figure 4: Visualization of annotated data sample from GenS-Video-150K.